MARS DESERT RESEARCH STATION

Sol Summary – March 17th

MDRS Daily Summary Report for SOL 6

Person filling out Report: Michał Kazaniecki, Engineer

Summary Title: It’s a good day to explore Mars

Mission Status: we can focus more on our field work instead of current hab maintenance

Sol Activity Summary: geo research during EVA, installing new pump, other work in hab with robot

Look Ahead Plan: EVA, mid-term summary session

Anomalies in work: our first day with every hab system working nearly properly

Weather: sunny, no strong wind

Crew Physical Status: one of crew members has mild cold sympthoms

Reports to be file:

-Operations report
-Commander report
-EVA Report
-EVA Request

Sol Summary – March 16th

MDRS Daily Summary Report for SOL 5

Person filling out Report: Michał Kazaniecki, Engineer

Summary Title:  Team is finally complete

Mission Status: successful EVA, progress with current work

Sol Activity Summary: training for 5th crew member, EVA, maintenance, green hab work

Look Ahead Plan: further work with robot, EVA,

Anomalies in work: mysterious water leakage outside the hab getting even bigger, need to manually refill static tank without a pomp

Weather: sunny, winds in the evening

Crew Physical Status: everything is correct

Reports to be file:

-Operations report
-Commander report
-Green Hab Officer’s report
-EVA request

Sol Summary – March 13th

MDRS Daily Summary Report for SOL 2

Person filling out Report: Michał Kazaniecki, Engineer

Summary Title: Last day of freedom

Mission Status: Final prepearation for SIM

Sol Activity Summary: reconnaissance outside the hab (and surrounding roads), fixing the tunnel, geological survey

Look Ahead Plan: first day in fully isolation (SIM)

Anomalies in work: solar panels still don’t work properly

Weather: sunny, calm, no strong wind

Crew Physical Status: minor injury while opening the can, but everyone is still alive and in good condition

Reports to be file:

-Operations report

-Commander report

-Mission Plan

-Bios and Photos

-Science report

Sol Summary – March 2nd

SOL SUMMARY REPORT SOL 18
SOL: 18

Person filling out Report: Louis MALLER, XO

Summary Title: The Artists

Mission Status: successful silent EVA in the morning, all systems go, work on experiments ongoing

Sol Activity Summary: EVA, science work on different experiments, sport

Look Ahead Plan: Tomorrow we will go for a last EVA, debrief the mission, and prepare to exit the Sim.

Anomalies in work: gave up on the 3D printer, clock in the living quarters not functional

Weather: sunny, no wind

Crew Physical Status: Injured crewmember conditions stable, one crewmember with a slight headache, rest of crew feeling well

EVA: non-verbal EVA for balloon debris recovery, AR glasses tests and seismometer recovery – success

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #17 report
– EVA #18 request
– Science Report

Sol Summary – March 1st

SOL SUMMARY REPORT SOL 17
SOL: 17

Person filling out Report: Louis MALLER, XO
Summary Title: Up!
Mission Status: successful EVA in the morning, all systems go, work on experiments ongoing
Sol Activity Summary: EVA, science work on different experiments, sport
Look Ahead Plan: Tomorrow we will gather the seismometer, then explore some more as we prepare to leave
Anomalies in work: gave up on the 3D printer, clock in the living quarters not functional, balloon untethered and seems to have flown away
Weather: very sunny, some mild winds
Crew Physical Status: Injured crewmember conditions stable, rest of crew feeling well
EVA: EVA for balloon deployment, AR glass tests and some around the Hab exploration

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #16 report
– EVA #17 request
– Astronomy Report
– Science Report

Sol Summary – February 28th

SOL SUMMARY REPORT SOL 16
SOL: 16

Person filling out Report: Louis MALLER, XO

Summary Title: Rose-tinted glasses

Mission Status: successful EVA in the afternoon, all systems go, work on experiments ongoing

Sol Activity Summary: sport, EVA, science work on different experiments

Look Ahead Plan: Hoping to have good weather for 24-hour balloon deployment tomorrow

Anomalies in work: 3D printer being troubleshooted, clock in the living quarters not functional

Weather: very sunny, some mild winds

Crew Physical Status: Injured crewmember conditions stable, rest of crew feeling well

EVA: EVA for engineering check and some pictures for emergency protocols report

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #15 report
– EVA #16 request
– Science Report

Sol Summary – February 27th

SOL SUMMARY REPORT SOL 15
SOL: 15
Person filling out Report: Louis MALLER, XO
Summary Title: Betelgeuse Broadcasting Corporation
Mission Status: successful EVA in the late morning, all systems go, work on experiments ongoing
Sol Activity Summary: brunch, EVA, science work on different experiments
Look Ahead Plan: Hoping to have good weather for 24-hour balloon deployment
Anomalies in work: 3D printer being troubleshooted, clock on the first floor not working correctly (little arm stuck between 6 and 7 O’clock)
Weather: mostly sunny, very strong gusts of winds.
Crew Physical Status: Injured crewmember feeling better, rest of crew feeling well
EVA: EVA with visiting journalists from the Betelgeuse Broadcasting Corporation, sent by Satellite 5, checked the seismometer, tested the alternative Martian navigation system, and explored moon- and mars- like regions.

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #14 report
– EVA #15 request
– HSO report
– Astronomy Report
– Science Report

Sol Summary – February 26th

SOL SUMMARY REPORT SOL 14
SOL: 14
Person filling out Report: Louis MALLER, XO
Summary Title: Brunch on Mars
Mission Status: successful EVA in the afternoon, all systems go, work on experiments ongoing
Sol Activity Summary: brunch, EVA, science work on different experiments
Look Ahead Plan: Looking forwards to Brian Cox joining us tomorrow for a Sol on Mars.
Anomalies in work: 3D printer being troubleshooted
Weather: beautiful and sunny, breeze with gusts of winds.
Crew Physical Status: Injured crewmember feeling better, rest of crew feeling well
EVA: EVA in front of the Hab, in order to learn, test and refine security procedures, went quite well.

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #13 report
– EVA #14 request
– HSO report
– Astronomy Report

Sol Summary – February 25th

SOL SUMMARY REPORT SOL 13

SOL: 13
Person filling out Report: Louis MALLER, XO
Summary Title: Exploring Cold Mars
Mission Status: successful EVA in the morning, all systems go, work on experiments ongoing
Sol Activity Summary: sport, EVA, science work on different experiments
Look Ahead Plan: Tomorrow there should be an EVA aimed at testing the feasibility of safety protocols designed today
Anomalies in work: 3D printer being troubleshooted, pumps frozen in the morning (thawed later in the morning)
Weather: beautiful and sunny in the morning slight breeze, quite cold
Crew Physical Status: One crewmember fell from the ladder to the loft, bruised but nothing broken
EVA: tried to fill in the static tank at the beginning of the EVA, failed because the pump had frozen. Succesfully deployed the balloon in the plain north of the Hab. Went to Explore the Moon again, and then headed southwards towards Skyline Rim. Then we headed back to the Hab, folded up the balloon, and managed to activate the pump.

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #12 report
– EVA #13 request
– HSO report
– Astronomy Report

Sol Summary – February 24th

SOL SUMMARY REPORT SOL 12
Person filling out Report: Louis MALLER, XO
Summary Title: Outlandish landscapes!

Mission Status: successful EVA in the morning, all systems go, work on experiments ongoing
Sol Activity Summary: sport, EVA, science work on different experiments
Look Ahead Plan: Tomorrow there should be an EVA, hoping for less windy weather for balloon deployment, and other experiments can also be tested outside.
Anomalies in work: 3D printer being troubleshooted
Weather: beautiful and sunny in the morning, but quite cold and with irregular blasts of wind
Crew Physical Status: Crew feeling well

EVA: tried to deploy the balloon on the location we had identified north of the hab, but failed because of heavy and sudden gusts of wind. Then we headed north, scouted for interesting places, took some pictures, and were particularly stunned by the place aptly named “The Moon” on the MDRS Map. Truly out-of-this world. We felt like Apollo astronauts as our vehicles and boots lifted up fine gray dust. Weather cold and windy but blue sky and absolutely no fogging in helmets, so we could fully enjoy the wonders of our Mars.

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #11 report
– EVA #12 request
– Science report
– 3D Printer status report

3D Printer Status
We have set to the task of making the 3D printer work, as we had some free time on our hands. We found the task more difficult than we had previously thought. Indeed, last year, when Arthur and I were on the station the 3D printer was perfectly working. It had no problem calibrating, and all the material was inside the box (with the possible exception of the Gap Gauge, sandpaper and Allan Wrench, citing by memory as this was one year ago).. It had a USB key in which we had put the executable for the printer and a few printable files.

When we opened the box again today, we found the printer in a sorry state. The scraping spatula, the quick start guide, the pliers, the casing and its content (including USB key)… had disappeared.

The cube glue is almost empty. There is only one cartridge left (which seems logical as it the other one has probably been used up), but it wasn’t operational.

We first decided to repair the cartridge, as we found it in three different pieces which is not its normal state. We then tried to calibrate the printer. We think that one of the three knobs that are supposed to maintain the plate is broken. Unfortunately, we didn’t manage to repair it and we couldn’t calibrate the plate so our test print failed. One of the issues might be a malfunction of the printing head as the melted plastic does not seem to come out of it and as the color of the plastic is mostly grey even if it is a black colored cartridge. This is the next step we will look forward to fix the printer but it’s an issue we still have to work on. We also hope to fix the knob maybe directly by using glue.